(*) If x = tan A – tan B, y = cot B – cot A then 1/x + 1/y =

(*)
If x = tan A – tan B, y = cot B – cot A then 1/x + 1/y =

Grade:11

2 Answers

Y RAJYALAKSHMI
9 years ago
x = tan A – tan B = sin A / cos A – sin B / cosB
=(sinA cosB – cosA sinB)/ cosA cosB = sin(A – B)/cosA cosB
y = cotB – cotA = cosB/sinB – cosA/sinA
=(sinA cosB – cosA sinB)/sinA sinB = sin (A – B)/sinA sinB
1/x + 1/y = [cosA cosB/sin(A – B)] + [sinA sinB/ sin(A – B)]
= (cosA coB + sinA sinB )/sin(A – B) = cos (A – B) / sin(A – B) 
= tan (A – B)
 
Ans: tan(A – B)
 
Y RAJYALAKSHMI
9 years ago
Pl. change the answer to cot (A – B)